OpenStudy (anonymous):

Given two points on the line, input the equation of the line in standard form Ax + By = C. Reduce all fractional answers to lowest terms. (7, -3), (4, -8)

myininaya (myininaya):

Well we can put into y=mx+b form first then rearrange it to the standard form. So m represents slope and b is the y-intercept. Given two points, you can use the slope formula which is (x1-x2)/(y1-y2)

OpenStudy (anonymous):

So (7-4)/(-3-(-8)?

myininaya (myininaya):

oops i put my x's on top. (y1-y2)/(x1-x2)

OpenStudy (anonymous):

Ok so (-3-(-8)/(7-4)?

myininaya (myininaya):

yes (-3-(-8))/(7-4) Do the subtraction on top and then do the subtraction on bottom.

OpenStudy (anonymous):

So 5/3

myininaya (myininaya):

yes that is m, the slope so let's fill in the info we have y=5x/3+b The only thing we don't know now is b, the y-intercept. you can use a point on the line (either of the two given to you) to find b. So if we use (7,-3) then we will see we have only one know in the following equation -3=5(7)/3+b Solve for b.

OpenStudy (anonymous):

So -3*3=5(7)/3*3+b = 9=35+b?

myininaya (myininaya):

Well if you are multiplying both sides by 3 to get rid of that fraction then you can't forget to multiply each term on the right hand side by 3. So you should have -3*3=5(7)/3*3+3*b

myininaya (myininaya):

-9=35+3b

OpenStudy (anonymous):

Ok. So next I need to get 3b on the left side right?

myininaya (myininaya):

Well you could leave it on the right and do something with that 35 instead but yeah either way really.

OpenStudy (anonymous):

Okay so -9 - 35 = 35 - 35 +3b?

myininaya (myininaya):

yep yep

OpenStudy (anonymous):

Okay so -44 = 3b?

myininaya (myininaya):

yep

OpenStudy (anonymous):

And then -44/3 = 3b/3 which equals -14 2/3 =3b?

myininaya (myininaya):

Yes I would leave it just in the improper form. So we have m=5/3 and b=-44/3 This means y=mx+b ---> y=5x/3-44/3

myininaya (myininaya):

We should be able to rearrange this into the form you want Ax+By=C

OpenStudy (anonymous):

So 5x/3+y=-44/3?

myininaya (myininaya):

Did you subtract 5x/3 on both sides? Looks like you subtracted on one side but not the other.

OpenStudy (anonymous):

So -5x/3+y=-44/3?

myininaya (myininaya):

Yes. Now get rid of fractional part by multiplying both sides by 3 and you are done.

OpenStudy (anonymous):

So -5x + y = -44?

myininaya (myininaya):

You need to also multiply that one term by 3.

myininaya (myininaya):

3(-5x/3+y=-44/3) 3(-5x/3+y)=3(-44/3) 3(-5x/3)+3(y)=3(-44/3) -5x + 3y = -44

OpenStudy (anonymous):

Ohh! Alright! Thank you!

myininaya (myininaya):

You did good. Good job Dani.